OpenStudy (anonymous):

let f(x)=9x-2 and g(x)=-x+3 find f(g(x))

OpenStudy (anonymous):

\[ g(x) = -x + 3 \] \[ f(x) = 9x - 2 \] \[ f(g(x)) = 9g(x) - 2 \] \[ f(g(x)) = 9(-x + 3) - 2 \] \[ f(g(x)) = -9x + 27 - 2 \] \[ f(g(x)) = -9x + 25 \]